Circular No. 5994

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SCHWASSMANN-WACHMANN 1 E. Lellouch, Observatoire de Paris, Meudon (OPM); G. Paubert, In-

stitut de Radio Astronomie Millimétrique (IRAM); N. Biver, D. Bockelée-Morvan, P. Colom, J. Crovisier, and L. Jorda, OPM; P. Rocher, Bureau des Longitudes, Paris; and D. Despois, Observatoire de Bordeaux, report: “We observed the CO J(2–1) line in P/Schwassmann-Wachmann 1 at the IRAM 30-m telescope on May 10–11 with a 0.06-km/s spectral resolution. The line consisted of a very sharp component (0.15 km/s wide) of 0.6 K in brightness temperature, blueshifted by −0.5 km/s, superimposed on a component of ∼ 0.1 K extending from −0.9 to +0.5 km/s. A contamination of the broader feature by galactic CO emission gradients cannot be excluded at the present time. The blueshift of the sharp peak is indicative of outgassing from the illuminated side of the nucleus, but its value is larger than expected from the sublimation of exposed CO ice. The total line area was 0.19 ± 0.02 K km s^{−1}, which agrees well with the line area previously observed at the JCMT (IAUC 5929), taking into account beam sizes and efficiencies. The corresponding production rate was of the order of 2 × 10^{28} molecules/s (or 10^{28} when taking into account only the sharp component).”

4U 0142+614 AND RX J0146.9+6121 C. Hellier, Department of Astronomy, University of Texas at Austin,

communicates: “A 10 700-s ROSAT PSPC pointed observation on 1993 Feb. 12–13 confirms that 4U 0142+614 is the source of an 8-s pulsation (cf. IAUC 5889). The heliocentric period is 8.6878 ± 0.0001 s and the semi-amplitude 5 percent, close to values from the 1984 EXOSAT ME data. The mean count rate was 4.8 counts/s. The PSPC also recorded RX J0146.9+6121 at a rate of 0.18 counts/s. It has a strong pulsation of 60 percent semiamplitude with a period of 1413 ± 8 s, confirming that this star (identified with the Be star LSI +61^{o}235) is the source of the pulsation previously attributed to 4U 0142+614, as suggested by Mereghetti et al. (1993, A.Ap. 278, L23). The period is significantly different from the value of 1455 ± 3 s in the 1984 EXOSAT data (White et al. 1987, MNRAS 226, 645). This is the first time that the two periodicities have been securely assigned to their correct stars. RX J0146.9+6121 is probably a Be/neutron-star binary with the slowest known neutron-star rotation. The high L_{x}/L_{opt} of > 10 for 4U 0142+614 suggests that it is a low-mass x-ray binary.”

1994 May 24 Daniel W. E. Green